write in slope-intercept form the equation of the line that passes through the given point and has the given slope (4,-2), m=1/2
Answer by haileytucki(390)

(4,-2),m=(1)/(2)
Find the value of b using the formula for the equation of a line.
y=mx+b
Substitute the value of m into the equation.
y=((1)/(2))*x+b
Substitute the value of x into the equation.
y=((1)/(2))*(4)+b
Substitute the value of y into the equation.
(-2)=((1)/(2))*(4)+b
Since b is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
((1)/(2))*(4)+b=(-2)
Multiply ((1)/(2)) by (4) to get ((1)/(2))(4).
((1)/(2))(4)+b=(-2)
Remove the parentheses around the expression -2.
((1)/(2))(4)+b=-2
Multiply (1)/(2) by 4 to get 2.
(2)+b=-2
Reorder the polynomial 2+b alphabetically from left to right, starting with the highest order term.
b+2=-2
Find the value of b.
b=-4
Now that the values of m(slope) and b(y-intercept) are known, substitute them into y=mx+b to find the equation of the line.
y=(x)/(2)-4